# Machine-assisted classification of potential biosignatures in earth-like exoplanets using low signal-to-noise ratio transmission spectra

### v1

@@ 2.3 Multilabel classification
where $N$ is the total number of instances (i.e., planets), $L$ is the number of labels per instance (i.e., number of molecular species), and $I$ is the indicator function that returns 1 if the condition is true and 0 otherwise. The condition for this metric is that the $j$th label the algorithm associates with the $i$th instance, $\hat{y}_{ij}$, is not the correct one, $y_{ij}$. For instance, if the algorithm predicts that planet 1 has molecule 2 when it does not, then $I(y_{12}\neq\hat{y}_{12})=1$.
Hamming Loss measures the fraction of incorrectly predicted labels. For example, if we need to determine if a planet has three molecular species and the algorithm incorrectly assigns two of the three labels, the Hamming Loss is $0.\overline{6}$. A large Hamming Loss generally indicates worse multilabel classification performance.
@@
In summary, evaluating the performance of a classification algorithm is not a trivial problem. Depending on which aspects of the classification we want to focus different metrics can be used. In this work we will apply a combination of binary and multilabel classification metrics to assess the capabilities of our algorithms.
3 The case of TRAPPIST-1 and its planets
@@
In order to test our methods and techniques, we need to select a proper investigation case that lies between an idealized illustration-purposed planet and a true target. After considering several possibilities, we found that the case of TRAPPIST-1 e fulfills the conditions for our numerical experiments. In the following paragraphs, we will present a summary of the potential that the TRAPPIST-1 system in general and TRAPPIST-1 e, in particular, have in the search for biosignatures.
@@
The TRAPPIST-1 system has gained significant scientific attention in recent years, especially in planetary sciences and astrobiology, owing to its exceptional features. The star, with a spectral type of $M8.0\pm 0.5$ (Gillon et al., 2016), is known for having the highest number of rocky planets discovered. This makes it an ideal candidate for atmospheric detection, especially because, besides other stars within 40 pc and using JWST, we can achieve favorable signal-to-noise ratio levels ($\text{SNR}\sim 5.5$) in a not-unrealistic observing time (200 h) (Gillon et al., 2020).
@@
Despite being approximately 7.6 billion years old (Burgasser & Mamajek, 2017) and displaying moderate magnetic activity compared to other stars of its class and age (Gillon, 2024), the emission of XUV radiation remains significantly high, suggesting that it could potentially erode the atmospheres of the surrounding planets (Wheatley et al., 2017; Bourrier et al., 2017).
@@
The orbital configuration of the TRAPPIST-1 system suggests that they may have migrated from farther regions, potentially leading to a rich accumulation of volatiles (Huang & Ormel, 2022; Agol et al., 2021). This makes it an excellent laboratory to study rocky planets near ultracool dwarfs and their atmospheric retention capacity (Airapetian et al., 2020). This is important for understanding habitability around such stars, which may host most potentially habitable planets in our galaxy (Gillon, 2024).
@@
Research on atmospheric stability suggests that these planets underwent phases of runaway greenhouse effect and intense desiccation due to stellar radiation before the star reached the main sequence (Luger & Barnes, 2015). Currently, planets e, d, and f reside within the conservative habitable zone, implying a conducive environment for maintaining secondary atmospheres. Models indicate the possible existence of atmospheres dominated by CO2, such as those of Venus and Mars in the Solar System. Under more restricted conditions, their atmospheres can also be dominated by other molecular species, such as H2O, O2, N2, CH4, or NH3 (Turbet et al., 2020). The case of Earth is one of a planet dominated by N2.
@@
Regarding the composition of these exoplanets, measurements indicate they are primarily rocky, with TRAPPIST-1 e standing out for a density that makes it closest to Earth in comparative terms, though not identical in iron proportion or volatile content. These data allow for the proposition of liquid oceans based on the projection that TRAPPIST-1 e could have an iron fraction of 25% (Agol et al., 2021). The astrobiological interest in TRAPPIST-1 e is considerable, given its potential to present a wide range of atmospheric compositions, as the density suggests a high molecular mass atmosphere (Turbet et al., 2020; Lin et al., 2021), with the possibility of maintaining liquid surfaces under various atmospheric compositions and pressures (Turbet et al., 2018). This potential is enhanced by the estimation that the initial water loss, due to the runaway greenhouse effect, was significantly less compared to the inner planets of the system (Turbet et al., 2020).
@@
Research into detecting Earth-like atmospheres in TRAPPIST-1 e using the JWST has motivated a variety of studies focused on specific aspects. On one hand, simulations of high-resolution spectra aim to replicate the spectral signature of Earth as seen from TRAPPIST-1 e, providing a crucial theoretical framework for identifying biosignatures (Kaltenegger et al., 2020; Lin & Kaltenegger, 2022; Lustig-Yaeger et al., 2023). On the other hand, the capabilities of JWST instruments to detect specific biosignatures face significant challenges in detecting molecules like O3 and N2O, which would exceed the telescope’s operational time. However, other authors highlight the viability of identifying CH4, CO2 and H2O in a limited number of transits (Lin et al., 2021; Lustig-Yaeger et al., 2023). Additionally, research on the detection of stratospheric clouds concludes that, given the observational capabilities of the JWST, such detection is not yet feasible (Doshi et al., 2022).
@@
In the following section, we enumerate and describe the biosignatures we have selected for our numerical experiments with TRAPPIST-1 e. As mentioned before, some of the biosignatures we select here have been identified as the most problematic to detect using the present observational capabilities and for that reason require novel techniques as that proposed here.
4 Selected biosignatures
4.1 CH4 and CO2: primary and secondary biosignatures
@@
Methane is highly common in reducing atmospheres, such as hydrogen atmospheres, and can be an anaerobic product. On Earth, methane is mainly produced by bacteria and plays a crucial role in the greenhouse effect, though it also comes from non-biological sources (Schwieterman et al., 2018). This molecule is considered a biosignature, especially in oxidizing environments because oxidation processes with $OH$ constantly remove it (Grenfell, 2018) and its presence could only indicate a chemical disequilibrium created by a complex process such as life. For instance, Schwieterman et al. (2018) have proposed that methane would be a biosignature when O2 or O3 are present in an atmosphere since they are highly oxidizing molecular species. However, other molecules, such as CO2, which also reveal an oxidizing atmospheric state, could also support the case of the case of Methane as a biosignature.
@@
It should be noted that CH4 can also be continuously replenished on an uninhabited planet through the reaction of CO2 with liquid water via serpentinization, a process involving the hydration of ultramafic minerals such as olivine (Grenfell, 2018; Schwieterman et al., 2018). Therefore, while methane may not be a primary biosignature in this context, its presence still indicates the existence of liquid water and, consequently, of habitable conditions. Thus, methane can be considered a secondary biosignature in this scenario. Guzmán-Marmolejo et al. (2013) suggested that N2-CO2 dominated atmospheres with CH4 concentrations greater than 10 ppmv could only be produced via biological processes, providing a method to distinguish between abiotic and biotic sources of methane.
4.2 $O_{2}$ y $O_{3}$: The crown jewel
Molecular oxygen on Earth is primarily the result of oxygenic photosynthesis. It has been proposed that O2 is an excellent biosignature due to its scarce abiotic production under Earth-like conditions (Léger et al., 2011). However, other potential abiotic sources, such as the photolysis of H2O or CO2, depend on the host star, the state of the planet’s atmosphere, and internal processes that limit the reabsorption of O2 (Schwieterman et al., 2018). Even in the case of an abiotic source of O2, other molecules may help to disentangle its origin.
To distinguish between biotic and abiotic O2, one can look for specific spectral fingerprints. For instance, the simultaneous presence of CO and O2 can indicate CO2 photolysis, while strong O2-O2 CIA features can signify significant hydrogen loss and oxygen-enriched atmospheres. This hydrogen loss occurs through the photolysis of H2O, where the lighter hydrogen escapes to space, leaving behind oxygen that accumulates in the atmosphere (Schwieterman, 2016). The absence of N2 might also suggest abiotic processes due to the lack of a cold trap that retains N2 in the upper atmosphere, allowing it to escape more easily (Wordsworth & Pierrehumbert, 2014).
But detecting O2 in exoplanetary atmospheres can be challenging. Collision-induced absorption (CIA) by O2 molecules produces a spectral signature that, although detectable, requires a relatively close proximity to our solar system. In the case of the JWST, for instance, a maximum distance of 5 parsecs is required for achieving a successful retrieval (Fauchez et al., 2020).
@@ 4.2 $O_{2}$ y $O_{3}$: The crown jewel
On the other hand, O3 is formed from O2 via photochemical processes and therefore acts as a secondary biosignature, i.e. its detection suggests the presence of oxygen and hence biological activity. The concentration levels in an exoplanet atmosphere would depend heavily on the UV emission from the host star. In cool M-dwarfs, there is a Goldilocks zone for ozone concentration that varies depending on the peak of UV emission (Grenfell et al., 2014).
Ozone exhibits a strong spectral signature in the near and medium-infrared (NIR and MIR, respectively), making it more detectable than O2 (Schwieterman et al., 2018) at least for instruments like the JWST’s NIRSpec and MIRI.
@@
The simultaneous presence of O2 (detected indirectly with O3) and CH4 in an exoplanet would indicate an atmosphere in redox disequilibrium. This combination could suggest an atmosphere actively regulated by life, as these molecules react quickly to produce H2O and CO2 (Lin et al., 2021; Schwieterman et al., 2018).
@@
Even though many more molecular species have been proposed as biosignatures (see e.g. Grenfell 2018; Schwieterman et al. 2018 and references therein), in our numerical experiments, we will focus on CH4 and O3 to demonstrate the capabilities of our approach. In future works, we expect to widen the set of molecular species for studying the detectability of other biosignatures.

### v4

@@ 2.3 Multilabel classification
where $N$ is the total number of instances (i.e., planets), $L$ is the number of labels per instance (i.e., the number of molecular species), and $I$ is the indicator function that returns 1 if the condition is true and 0 otherwise. The condition for this metric is that the $j$th label the algorithm associates with the $i$th instance, $\hat{y}_{ij}$, is not the correct one, $y_{ij}$. For instance, if the algorithm predicts that planet 1 has molecule 2 when it does not, then $I(y_{12}\neq\hat{y}_{12})=1$.
Hamming Loss measures the fraction of incorrectly predicted labels. For example, if we need to determine whether a planet has three molecular species and the algorithm incorrectly assigns two of the three labels, the Hamming Loss is $0.\overline{6}$. A large Hamming Loss generally indicates worse multilabel classification performance.
@@
In summary, evaluating the performance of a classification algorithm is not a trivial matter. Depending on which aspects of the classification we want to focus on, different metrics can be used. In this work, we will apply a combination of binary and multilabel classification metrics to assess the capabilities of our algorithms.
3 The case of the star TRAPPIST-1 and its planets
@@
In order to test our methods and techniques, we need to select an appropriate investigation case that lies between an idealised, illustrative planet and a genuine target. After considering several possibilities, the TRAPPIST-1e system was identified as an ideal candidate for our numerical experiments. In the following paragraphs, we summarise the potential of the TRAPPIST-1 system in general, and TRAPPIST-1e in particular, in the search for biosignatures.
@@
The TRAPPIST-1 system has gained significant scientific attention in recent years, especially in planetary sciences and astrobiology, owing to its exceptional features. The star, with a spectral type of $M8.0\pm 0.5$ (Gillon et al., 2016), is known for hosting the highest number of rocky planets discovered so far. This makes it an ideal candidate for atmospheric detection, particularly as, compared to other stars within 40 pc, observations using JWST can achieve favourable signal-to-noise ratio levels ($\text{SNR}\sim 5.5$) within a realistic observing time (200 h) (Gillon et al., 2020).
@@
Despite being approximately 7.6 billion years old (Burgasser & Mamajek, 2017) and displaying moderate magnetic activity compared to other stars of its class and age (Gillon, 2024), its XUV radiation emission remains significantly high, suggesting that it could potentially erode the atmospheres of the surrounding planets (Wheatley et al., 2017; Bourrier et al., 2017).
@@
The orbital configuration of the TRAPPIST-1 system suggests that they may have migrated from farther regions, potentially leading to a rich accumulation of volatiles (Huang (HuangShuo) & Ormel, 2022; Agol et al., 2021). This makes it an excellent laboratory for studying rocky planets orbiting ultracool dwarfs and their atmospheric retention capacity (Airapetian et al., 2020). This is important for understanding habitability around such stars, which may host most potentially habitable planets in our galaxy (Gillon, 2024).
@@
Research on atmospheric stability suggests that these planets underwent phases of runaway greenhouse effect and intense desiccation due to stellar radiation before the star reached the main sequence (Luger & Barnes, 2015). Currently, planets e, d, and f reside within the conservative habitable zone, suggesting favorable conditions for retaining secondary atmospheres. Models indicate the possible existence of atmospheres dominated by CO2, such as those of Venus and Mars in the Solar System. Under more restricted conditions, their atmospheres can also be dominated by other molecular species, such as H2O, O2, N2, CH4, or NH3 (Turbet et al., 2020). The case of Earth is that of a planet dominated by N2.
@@
Regarding the composition of these exoplanets, measurements indicate that the planets are primarily rocky, with TRAPPIST-1 e standing out for a density that makes the system closest to Earth in comparative terms, though not identical in iron proportion or volatile content. These data support the potential existence of liquid-water oceans based on the projection that TRAPPIST-1 e could have an iron fraction of 25% (Agol et al., 2021). TRAPPIST-1 e is of significant astrobiological interest, given its potential to present a wide range of atmospheric compositions, as the density suggests a high molecular mass atmosphere, with the possibility of maintaining liquid surfaces under various atmospheric compositions and pressures (Turbet et al., 2020; Lin et al., 2021). This potential is enhanced by the estimation that the initial water loss, due to the runaway greenhouse effect, was significantly less than that of the inner planets of the system (Turbet et al., 2020).
@@
Research into detecting Earth-like atmospheres in TRAPPIST-1 e using the JWST has motivated a variety of studies focused on specific aspects. On one hand, simulations of high-resolution spectra aim to replicate the spectral signature of Earth as seen from TRAPPIST-1 e, providing a crucial theoretical framework for identifying biosignatures (Kaltenegger et al., 2020; Lin & Kaltenegger, 2022; Lustig-Yaeger et al., 2023). On the other hand, the capabilities of JWST instruments to detect specific biosignatures face significant challenges in detecting molecules like O3 and N2O, which would surpass feasible observational time allocations for the telescope. Nevertheless, other studies emphasize that detecting CH4, CO2, and H2O is viable within a limited number of transits (Lin et al., 2021; Lustig-Yaeger et al., 2023). Additionally, research on the detection of stratospheric clouds concludes that, given the observational capabilities of the JWST, such detection is not yet feasible (Doshi et al., 2022).
@@
In the following section, we enumerate and describe the biosignatures/bioindicators we have selected for our numerical experiments with TRAPPIST-1 e. As mentioned before, some of the biosignatures/bioindicators we select here have been identified as the most problematic to detect using the present observational capabilities and, for that reason, require novel techniques such as the one proposed here.
4 Selected biosignatures and bioindicators
4.1 CH4 and CO2: primary and secondary bioindicators
@@
Methane is common in reducing atmospheres, such as hydrogen-rich atmospheres, and may be produced under anaerobic conditions. On Earth, methane is mainly produced by bacteria and plays a crucial role in the greenhouse effect, although it also originates from non-biological sources (Schwieterman et al., 2018). This molecule is considered a bioindicator, especially in oxidising environments because oxidation processes involving $OH$ radicals continuously remove it, so its presence may indicate a chemical disequilibrium arising from complex processes such as life. For instance, Schwieterman et al. (2018) have proposed that methane serves as a biosignature when O2 or O3 are present in an atmosphere, since these are highly oxidising molecular species. However, other molecules, such as CO2, which also reflect an oxidising atmospheric state, could support the case for methane as a bioindicator.
@@
It should be noted that CH4 can also be continuously replenished on an uninhabited planet through the reaction of CO2 with liquid water via serpentinisation, a process involving the hydration of ultramafic minerals such as olivine (Grenfell, 2018; Schwieterman et al., 2018). Therefore, while methane may not be a primary biosignature/bioindicator in this context, its presence still indicates the existence of liquid water and, consequently, habitable conditions. Thus, methane can be considered a secondary biosignature in this scenario. Guzmán-Marmolejo et al. (2013) suggested that N2-CO2 dominated atmospheres with CH4 concentrations greater than 10 ppmv could only be produced via biological processes, providing a method to distinguish between abiotic and biotic sources of methane.
4.2 O2 and O3
Molecular oxygen (O2) on Earth is primarily the result of oxygenic photosynthesis. Historically, its detection on an exoplanet was considered a robust biosignature, as abiotic production under Earth-like conditions is minimal (Léger et al., 2011). However, recent studies have identified multiple abiotic pathways for O2 and O3 accumulation, prompting a reassessment of their reliability as indicators of life (Meadows et al., 2018; Schwieterman & Leung, 2024). These include photolysis of CO2, hydrogen escape, and other mechanisms dependent on planetary and stellar conditions.
One such abiotic mechanism is CO2 photolysis, which can generate O2 in dry, CO2-rich atmospheres. The recombination of CO and O into CO2 is spin forbidden, causing a slow reaction rate and enabling O2 accumulation, particularly under low humidity conditions (Gao et al., 2015). Additionally, on planets with surface oceans and humid atmospheres, UV radiation from M-dwarf stars can enhance CO2 photolysis, contributing to O2 and O3 accumulation (Domagal-Goldman et al., 2014; Harman et al., 2015). Another key scenario involves hydrogen loss in planets that have undergone a runaway greenhouse phase. During the pre-main-sequence stage of M-dwarf stars, their intense XUV irradiation can induce massive water loss, leading to preferential hydrogen escape and leaving behind oxygen (Luger & Barnes, 2015). Oceans play a critical role in this process, as extensive surface water reservoirs can provide a continuous source of H2O for photodissociation. The subsequent escape of hydrogen to space results in the accumulation of O2; depending on the efficiency of oxygen sinks (e.g., crustal oxidation, ion escape, reactions with volcanic gases), some planets may sustain oxygen-rich atmospheres without biological input (Schwieterman & Leung, 2024).
To differentiate biotic from abiotic O2, several spectral diagnostics can be employed:
The simultaneous presence of CO and O2 suggests CO2 photolysis, as CO would typically be removed in a biosphere with active oxidative processes (Schwieterman, 2016; Schwieterman & Leung, 2024).
O2–O2 Collision-Induced Absorption (CIA) at 1.06, 1.27, and 6.4 $\mu$m signals high O2 partial pressures, indicative of abiotic oxygen retention due to hydrogen loss (Fauchez et al., 2020).
A low abundance of N2 may suggest an atmosphere where the absence of non-condensable gases lifts the tropopause, enhancing H2O photodissociation and subsequent hydrogen escape (Wordsworth & Pierrehumbert, 2014).
@@ 4.2 O2 and O3
On the other hand, O3 is formed from O2 via photochemical processes and therefore acts as a secondary biosignature or bioindicator—its detection suggests the presence of oxygen and hence potential biological activity. However, models indicate that O3 levels are highly dependent on the UV output of the host star. In cool M-dwarfs, there is a Goldilocks zone for ozone concentration that depends on the peak of UV emission (Grenfell et al., 2014). This suggests that significant amounts of O3 could accumulate from lower O2 levels, which may have originated from abiotic processes.
O3 exhibits a strong spectral signature in the near-infrared (NIR) and mid-infrared (MIR), making it easier to detect than O2 with instruments such as JWST/NIRSpec and MIRI (Schwieterman & Leung, 2024). While the most prominent absorption peak occurs in the MIR range, Lustig-Yaeger et al. (2023) found that the 4.7 $\mu$m peak in the NIRSpec range allows for more precise retrievals. Despite the lower O3 signal at this wavelength, they highlight the importance of considering instrument sensitivity across different wavelengths, rather than solely relying on molecular absorption band strength when targeting specific molecules.
@@
The simultaneous presence of O2 (detected indirectly via O3) and CH4 in an exoplanet would indicate an atmosphere in redox disequilibrium (see, e.g., (Sagan et al., 1993)). This combination could suggest an atmosphere actively regulated by life, as these molecules react quickly to produce H2O and CO2 (Lin et al., 2021; Schwieterman et al., 2018).
@@
Even though many more molecular species have been proposed as biosignatures (see, e.g., Grenfell 2018; Schwieterman et al. 2018; Schwieterman & Leung 2024 and references therein), in our numerical experiments we will focus on CH4 and O3 to demonstrate the capabilities of our approach. In future works, we expect to widen the set of molecular species for studying the detectability of other potential biosignatures.
5 Mitigation of Stellar Contamination in Transmission Spectra
The accurate characterisation of exoplanetary atmospheres via transmission spectroscopy is significantly impacted by stellar contamination, particularly when observing planets orbiting active stars such as M dwarfs. Addressing this challenge requires a detailed understanding of the stellar contamination phenomenon and robust computational methods to effectively remove or mitigate its effects from observed spectra. In this context, we first describe the phenomenon of stellar contamination and subsequently introduce a machine learning technique employing denoising autoencoders specifically tailored to tackle these distortions.
5.1 Stellar contamination
One of the most significant obstacles when dealing with transmission planetary spectra, especially around relatively active stars, is the so-called stellar contamination. The effect, which is technically called transit light source effect (TLS), encompasses the alterations in the transmission spectra arising from heterogeneities in the star’s photosphere (Rackham et al., 2018, 2019, 2023). These heterogeneities are not easily revealed in photometric observations of the star and must often be modelled on a case-by-case basis (Rackham & De Wit, 2024).
Stellar heterogeneities, including cool starspots and hot faculae, differ from the average photosphere in both temperature and spectral output. Consequently, the light recorded during a transit includes not only the planet’s atmospheric signature but also uneven contributions from active, unocculted stellar regions. In Figure 2 we show several examples of how an original transmission spectra (first column in figure) is modified by the presence of stellar contamination (second column). As observed, stellar contamination may potentially mimic or mask genuine atmospheric features. Thus, for instance, these irregularities can hinder the detection of key molecules, such as water, especially in low-resolution observations (Genest et al., 2022). Moreover, stellar contamination can alter inferred planetary radii, inflate molecular abundance estimates, or reduce important atmospheric signals (Rackham et al., 2018).
The effect of stellar contamination is particularly critical for rocky exoplanets orbiting M dwarfs, such as TRAPPIST-1 e, which is our model planet here. These stars normally exhibit high magnetic activity and hence notable photospheric heterogeneities (Iyer & Line, 2020). In fact, stellar contamination has already been identified in transit observations conducted with JWST/NIRISS on TRAPPIST-1 b (Lim et al., 2023), confirming theoretical expectations.
Mitigation strategies involve employing models to correct false contamination signals or explicitly incorporating these effects into atmospheric retrieval schemes (Rackham & De Wit, 2024). Another approach uses out-of-transit stellar spectra to constrain surface heterogeneities before correcting the resulting transmission spectrum. However, these methods face substantial challenges. Current models often fail to accurately reproduce the spectra of M dwarfs, such as TRAPPIST-1, impeding effective contamination corrections (Rackham & De Wit, 2024). Furthermore, there is considerable degeneracy between the properties of stellar heterogeneities and planetary atmospheric characteristics, making it difficult to disentangle these contributions (Iyer & Line, 2020). Nowadays, more accurate models and additional observations are essential to better characterise stellar inhomogeneities and their effects on transmission spectra.
For the purpose of our numerical experiments, we have devised a strategy to remove stellar contamination from our synthetic spectra using the so-called de-noising autoencoders (see below).
Contaminated transmission spectra $\left(D_{\lambda,\text{obs}}\right)$ can be mathematically expressed as (Rackham et al., 2018):
$D_{\lambda,\text{obs}}=\epsilon_{\lambda}(f_{\mathrm{spot}},f_{\mathrm{fac}},T_{\mathrm{spot}},T_{\mathrm{fac}})\,D_{\lambda},$ (7)
where $D_{\lambda}$ is the nominal transit depth attributable solely to the planet, and the function $\epsilon_{\lambda}$ encodes the stellar contamination spectrum. The function $\epsilon_{\lambda}$ depends on the properties of the stellar surface heterogeneities, namely the spots and faculae. In particular, $f_{\mathrm{spot}}$ and $f_{\mathrm{fac}}$ denote the fraction of the stellar surface covered by spots and faculae, respectively, while $T_{\mathrm{spot}}$ and $T_{\mathrm{fac}}$ represent their respective temperatures. Moreover, the terms $F_{\lambda,\text{spot}}$, $F_{\lambda,\text{fac}}$, and $F_{\lambda,\text{phot}}$ correspond to the spectral flux densities at wavelength $\lambda$ for the spots, faculae, and the photosphere, respectively. According to Rackham et al., 2018 (Equation 3), a simplified expression for this contamination function is given by:
$\epsilon_{\lambda,\text{het}}=\frac{1}{1-f_{\text{spot}}\left(1-\frac{F_{\lambda,\text{spot}}}{F_{\lambda,\text{phot}}}\right)-f_{\text{fac}}\left(1-\frac{F_{\lambda,\text{fac}}}{F_{\lambda,\text{phot}}}\right)}.$ (8)
5.2 Autoencoder for denoising contaminated spectra
Autoencoders (AE) are versatile neural networks widely used in unsupervised learning. These neural networks are used for many different tasks including, but not restricted to, dimensionality reduction, anomaly detection, and noise removal (see, e.g., Berahmand et al. 2024 and references therein).
Typically, an autoencoder consists of two main components: an encoder network that compresses the input data into a so-called reduced latent space. In spectral analysis terms, the encoder converts a spectrum with, for instance, 100 channels or features into a smaller set (e.g., 10 features) that encodes the essential information from the original spectrum. The other component, the decoder, reconstructs the data from the compressed representation. The encoder and decoder pair are trained to minimise the loss produced during the encoding-decoding process.
Among all their variants, denoising autoencoders are specifically designed to handle noisy or corrupted inputs. During training, the model receives intentionally corrupted data as input and is tasked with reconstructing the original, uncorrupted data. This process effectively isolates and removes noise while preserving essential features (see, e.g., Berahmand et al. 2024).
In this study, we designed and developed a denoising autoencoder with a custom architecture tailored to mitigate stellar contamination and noise in transmission spectra. The model is implemented in Keras (Chollet, 2015) with a TensorFlow backend (Abadi et al., 2015). Our design incorporates multiple hidden layers with Swish activation functions, dropout regularisation to reduce overfitting, and a bottleneck layer for dimensionality reduction, preserving only the most salient features. The details of our design are beyond the scope of this paper and will be presented in a forthcoming publication. Nonetheless, all source code is available alongside the programs and data accompanying this work (see Data availability and reproducibility).
In Figure 2 we show several examples of applying our trained autoencoder to three representative synthetic spectra used in our numerical experiments. We can see how well our autoencoder denoises the contaminated and noisy spectra (panels in the second column in the figure), yielding clean versions (third column) that closely match the original uncontaminated spectra (first column).
The two-step approach—denoising and classification (see the next sections)—described in this paper underscores the promise of integrating state-of-the-art machine learning techniques, such as de-noising autoencoders, with established classification methods to tackle complex astrophysical problems.
Figure 2: Examples of the autoencoder implementation designed to mitigate stellar contamination in three representative synthetic transmission spectra: an airless planet (first row), a planet with a CO2-rich atmosphere (second row), and a planet with potential biosignatures (third row). The input spectra have an SNR of 3 and include stellar contamination levels of $f_{\mathrm{spot}}=0.08$ and $f_{\mathrm{fac}}=0.54$. The panels in the first column display the synthetic spectra free from noise and contamination; the second column shows the input spectra affected by noise and contamination; and the third column presents the reconstructed spectra returned by the autoencoder.